# High-inclination Centaur reservoirs beyond Neptune

### v1

@@ 3 Simulation method
Figure 2: TNO reservoirs at 4.54.5\,Gyr for different Tisserand parameters $T$. Hot TNOs are shown in red and cold TNOs are shown in blue. Unstable TNOs are shown in black at the last sampling epoch before instability. The first and third columns display the orbital distributions in the $(a,I)$ and $(a,e)$ planes. The second and fourth columns show II, $e$ and $a$ at minimum semimajor axis where a vertical black line denotes Neptune’s semi-major axis. The solid orange and blue curves in the inclination panels are respectively the Tisserand inclination pathway $I_{\rm Tiss}(a,T,a_{\rm N})$ (2) and the circular orbits’ initial inclination $I_{\rm circ}(a,T,a_{\rm N})$ (5) with respect to Neptune. The solid green curves for retrograde and prograde inclinations respectively are the Tisserand inclinations pathways with respect to Saturn $I_{\rm Tiss}(a,-2,a_{\rm S})$ ($135^{\circ}$ inclination) and $I_{\rm Tiss}(a,2,a_{\rm S})$ ($45^{\circ}$ inclination). In the eccentricity panels, the curves indicate perihelia that match the four giant planets’ semi-major axes. The Tisserand inclination pathway with respect to Neptune is not shown for $T=2.8$ because the expression (2) is not valid for $T>2.7$ (see Paper III for details).
Dynamical evolution is simulated 4.5 Gyr forward in time using the IAS15 $N$-body integrator of the REBOUND package (Everhart85; Rein15; Rein17). The simulations include the four giant planets, the TNO ensembles and passing stars. The terrestrial planets’ mass was added to the Sun’s. The three-dimensional Galactic tide (Heisler86) and relative inclination of the ecliptic and Galactic planes were included using the REBOUNDx package (Tamayo20). The Oort constants, $A=15.3$ km  s-1 kpc-1, $B=11.9$ km  s-1 kpc-1 , and star density in the solar neighbourhood, $\rho_{0}=0.119M_{\odot}$  pc-3, were obtained from Gaia’s first data release (Bovy17; Widmark19) and were used in Papers I and II.
@@
The orbits of 142 500 TNOs were integrated and their minimum semi-major axes monitored throughout the simulation. TNOs are removed from the simulation in case of collision or ejection.
@@
Figure 4: Average fraction of cold TNOs (in blue) and hot TNOs (in red) that survived 4.54.5\,Gyr as a function of the Tisserand parameter $T$ (and $I_{\infty}$).
@@
The end states of the TNO reservoirs in the $Ia$- and $ea$-planes are shown in Figure 2 for the Tisserand parameters, T=2,1.5,1, 0, 0.5, 2, 2.5T=-2,\ -1.5,\ -1,\ 0,\ 0.5,\ 2,\ 2.5 and $2.8$ (as in Figure 1). Hot TNOs are shown in red and cold TNOs in blue. Unstable orbits are shown in black at the last sampled epoch before instability. Also shown are the TNOs’ distributions of minimum semi-major axes achieved over 4.5 Gyr and the corresponding eccentricities and inclinations for all orbital types stable and unstable.
@@
The reservoirs’ semi-major axis extent is similar to the initial state’s. TNOs scattered by the combined effects of the giant planets, Galactic tide and passing stars are also present far outside the reservoirs’ domain. Secular perturbations cause a larger dispersion in eccentricity and inclination for hot TNOs. Cold TNOs remain mainly with low eccentricity orbits as evidenced by their clustering along the inclination curve of circular orbits (5). As the Tisserand parameter increases toward the polar corridor boundary’s $T=2$, instability is greater and the reservoirs are depleted. Examination of the minimum semi-major axis distributions shows that starting from T=2T=-2 minimum perihelia are distributed between Saturn and Neptune but the perihelion range narrows around Neptune’s orbit outside the polar corridor as $T$ increases and initial inclinations get smaller.
@@
The Tisserand inclination pathways with respect to Neptune for the bulk of the reservoirs’ material are generally followed for all values of the Tisserand parameter. The pathways are shown in orange in the inclination panels of Figure 2. A novel process related to Saturn adds secondary inclination pathways that are present for $T\leq 2$ and prominent inside the polar corridor. Saturn appears to retain a small fraction of reservoir material with perihelia at Saturn’s orbit whereas their orbits are pushed back outside Neptune’s. This material follows two specific Tisserand inclination pathways with respect to Saturn regardless of the initial Tisserand parameter, $T$. These are $I_{\rm Tiss}(a,-2,a_{\rm S})$ and $I_{\rm Tiss}(a,2,a_{\rm S})$ (2) where $a_{\rm S}$ is Saturn’s semi-major axis. The pathways correspond to inclinations far from Neptune $I_{\infty}=135^{\circ}$ and $I_{\infty}=45^{\circ}$ (3) and appear as the green curves in the inclination panels of Figure 2. These secondary populations are prominent for $T=0$ and $T=0.5$ in Figure 2 and their density decreases outside the polar corridor as $T$ tends to 2-2. Investigating Saturn’s secondary inclination pathways is beyond the scope of this paper.
@@
Figure 3 shows the ratio of the number of surviving hot TNOs to their initial number, denoted $f_{\rm TNO}$, for different initial semi-major axes as a function of the Tisserand parameter. For each semi-major axis, all the different perihelia are included in the statistic – the perihelion dependence is discussed in the next section. Four semi-major axis groups are visible. First, TNOs with initial $a$ in the range [70:100] au follow a similar distribution to that of the dynamical time in the three-body problem found in Paper IV (Figure 7 therein). The dynamical time may be considered a proxy for $f_{\rm TNO}$. The three-body dynamical time versus Tisserand parameter, has three peaks: a prominent one near $T\sim-0.7$, a moderate one near $T\sim 0.8$ and a smaller one near $T\sim 2.2$. The $6$-body plus passing stars configuration of the present framework displaces the three peaks to $T\sim-1.5$ with $f_{\rm TNO}\sim 60\%$ for $a=80$ au, to $T\sim 0.5$ with $f_{\rm TNO}\sim 50\%$ for $a=70$ au and to $T\sim 2$ with $f_{\rm TNO}\sim 30\%$ for $a=100$ au. This group’s range of [70:100] au is reminiscent of the semi-major axis probability distributions peaks discussed in Sections 1 and 2.
@@
The ensembles with $a=50$ au and $a=60$ au form the second group whose retrograde peak has been all but erased in favor a prominent peak at $T=0.5$ and $f_{\rm TNO}\sim 60\%$ for $a=60$ au. The third group has a semi-major axis range [110:140] au whose $f_{\rm TNO}$ decrease steadily with increasing $T$. The fourth group is made up of the ensembles of $a=40$ au, $a=200$ au and 500 au that are mainly unstable confirming the existence of two stability boundaries, an inner one between $40$ au and $50$ au and an outer one between $140$ au and $200$ au.
@@
The values of $f_{\rm TNO}$ in the range [50:140] au confirm that TNO reservoirs are long-lived in the presence of all giant planets, passing stars and the Galactic tide. In particular, the peaks at $I_{\infty}=122^{\circ}$ (T=1.5T=-1.5) and $I_{\infty}=80^{\circ}$ ($T=0.5$) single out inclinations that have the largest current populations. The average of $f_{\rm TNO}$ over all semi-major axes for hot and cold TNOs, denoted $\langle f_{\rm TNO}\rangle$, is shown a function of the Tisserand parameter in Figure 4. It indicates that cold TNOs on average have a 4.5 Gyr dynamical time as $\langle f_{\rm TNO}\rangle>50\%$ unsurprisingly since they mainly retain low eccentricity orbits. Hot TNOs fractions decrease with $T$ peaking at $T\sim-1.5$ with $\langle f_{\rm TNO}\rangle\sim 40\%$, $T\sim 0.5$ with $\langle f_{\rm TNO}\rangle\sim 35\%$ and $T\sim 2$ with $\langle f_{\rm TNO}\rangle\sim 22\%$.
@@ 5 Centaur injection
Figure 5: Fraction of hot TNOs that became Centaurs as a function of the Tisserand parameter $T$ (and $I_{\infty}$). In the top panel, $f_{\rm cent}$, corresponds to the initial semi-major axis given next to each curve. In the bottom panel, the semi-major axis-averaged, $\langle f_{\rm cent}\rangle$, of initially cold and hot TNOs are shown in blue and red respectively. The black line is $\langle f_{\rm cent}\rangle(\%)=17+4.3T$.
The ratio of the number of TNOs whose semi-major axes became smaller than Neptune’s over 4.5 Gyr, to the initial TNO number is denoted $f_{\rm cent}(a,T)$. The initial TNO number of each semi-major axis includes the full perihelion range and the Centaur number includes stable and unstable orbits. The top panel of Figure 5 shows $f_{\rm cent}(a,T)$ for hot TNOs and different semi-major axes as a function of the Tisserand parameter. The Centaur fraction increases with the Tisserand parameter at a larger rate as the semi-major axis is closer to Neptune. TNOs with low to moderate inclination are injected in greater numbers than high-inclination TNOs as their corresponding reservoirs are depleted because of instability. The hot TNO ensembles with $a=200$ au and 500 au contribute smaller Centaur numbers also because of instability. On the other hand, instability for TNOs with $a=40$ au translates into the largest injection rate because of its proximity to Neptune. For Tisserand parameters T<1T<-1, Centaur fraction is finite indicating that the presence of the giant planets and passing stars breaks Neptune’s reflecting ability that was evidenced in the three-body problem with the existence of $a_{\rm refl}$ (4) (Papers III and IV) – further evidence is given in the next section.
We define $\langle f_{\rm cent}\rangle$ as the semi-major axis averaged Centaur fraction in the stable region. The TNO ensembles of $a=40$ au, 200 au and 500 au are excluded from the hot TNO statistics because they lie outside the stability region. This will be applied in the remainder of the paper when averaging over semi-major axes. The bottom panel of Figure 5 shows that $\langle f_{\rm cent}\rangle$ of hot and cold TNOs are remarkably similar. The hot TNO $\langle f_{\rm cent}\rangle$ follows a simple linear expression that depends on the Tisserand parameter as follows $\langle f_{\rm cent}\rangle(\%)=17+4.3T$.
@@
More information on the injection process is found by examining the dependence of the Centaur number on perihelion for hot TNOs and on semi-major axis for cold TNOs. Figure 6 shows non-averaged $f_{\rm cent}(a,T,q)$ along with the corresponding dynamical time, $\tau_{d}(a,T,q)$, for the ensembles with an initial $a=80$ au. As stated in the Section 2, each $(a,T,q)$ ensemble has 100 TNOs and $f_{\rm cent}(a,T,q)$ is the fraction of those among them that became Centaurs. An ensemble’s dynamical time is its median lifetime. It is meant as a stability indicator and not a dynamical time in the classical sense which is associated with a single orbit and derived from its clones’ evolution. As such the TNOs’ dynamical time, $\tau_{d}$, shown here is an average of the classical dynamical times of the orbits that make up the $(a,T,q)$ ensemble. It is found unsurprisingly that Centaur injection tends to increase with smaller perihelia for low to moderate inclinations ($T=2.8$). However, the perihelion dependence become shallower for reservoirs inside the polar corridor ($1\leq T\leq 2$) suggesting that it is secular perturbations and not the initial perihelion that control Centaur injection – see the panel of $T=1$. The dynamical time is an increasing function of perihelion and reaches 4.5 Gyr in the polar corridor and for T>1T>-1. Its dependence on the Tisserand parameter inside the polar corridor is unclear but could be depend on the vicinity of $T$ to the boundaries T=1T=-1 and $T=2$. Different initial semi-major axes of hot TNOs in the region [50:140] au have similar $f_{\rm cent}(a,T,q)$ and dynamical times as a function of perihelion. The median and maximum dynamical times for the different perihelia respectively are 3 Gyr and 4.5 Gyr with the latter achieved in the polar corridor. The median and maximum dynamical times are 2.4 Gyr and 3.5 Gyr for $a=200$ au and 500 au, whereas for $a=40$ au, they are 0.48 Gyr and 1.5 Gyr.
Figure 7: Dynamical time (green) and Centaur fraction (purple) for initially cold TNOs as a function of semi-major axis for different Tisserand parameters. The dashed blue line is $\tau_{d}=4.5$ Gyr. The empty semi-major axis ranges indicate absence of Centaur injection.
The Centaur fraction, $f_{\rm cent}(a,T,q)$, and dynamical time of cold TNOs, $\tau_{d}(a,T,q)$, are shown in Figure 7. The number of injected Centaurs decreases exponentially with semi-major axis so that most injected Centaurs come for the smallest semi-major axes. This is visible on the dynamical times as they reach 4.5 Gyr as soon as $a\geq 60$ au regardless of the Tisserand parameter. It is interesting that regardless of which semi-major axis and perihelion contributes most to Centaur injection, the average Centaur fractions, $\langle f_{\rm cent}\rangle$, of both hot and cold reservoirs are similar (Figure 5 bottom panel). Understanding this similarity’s origin may uncover the common dynamical process underlying Centaur injection.

### v2

@@ 3 Simulation method
Figure 2: TNO reservoirs at 4.54.5\,Gyr for different Tisserand parameters $(T$). Hot TNOs are shown in red and cold TNOs in blue. Unstable TNOs are shown in black at the last sampling epoch before instability. The first and third columns display the orbital distributions in the $(a,I)$ and $(a,e)$ planes. The second and fourth columns show II, $e,$ and $a$ at minimum semimajor axis, where a vertical black line denotes Neptune’s semimajor axis. The solid orange and blue curves in the inclination panels are, respectively, the Tisserand inclination pathway $I_{\rm Tiss}(a,T,a_{\rm N})$ (2) and the circular orbits’ initial inclination, $I_{\rm circ}(a,T,a_{\rm N})$ (5), with respect to Neptune. The solid green curves for retrograde and prograde inclinations respectively are the Tisserand inclinations pathways with respect to Saturn $I_{\rm Tiss}(a,-2,a_{\rm S})$ ($135^{\circ}$ inclination) and $I_{\rm Tiss}(a,2,a_{\rm S})$ ($45^{\circ}$ inclination). In the eccentricity panels, the curves indicate perihelia that match the four giant planets’ semimajor axes. The Tisserand inclination pathway with respect to Neptune is not shown for $T=2.8$ because expression (2) is not valid for $T>2.7$ (see Paper III for details).
Dynamical evolution is simulated 4.5 Gyr forward in time using the IAS15 $N$-body integrator of the REBOUND package (Everhart85; Rein15; Rein17). The simulations include the four giant planets, the TNO ensembles and passing stars. The terrestrial planets’ mass was added to the Sun’s. The three-dimensional Galactic tide (Heisler86) and relative inclination of the ecliptic and Galactic planes were included using the REBOUNDx package (Tamayo20). The Oort constants, $A=15.3$ km  s-1 kpc-1, $B=11.9$ km  s-1 kpc-1 , and star density in the solar neighborhood, $\rho_{0}=0.119M_{\odot}$  pc-3, were obtained from Gaia’s first data release (Bovy17; Widmark19) and were used in Papers I and II.
@@
The orbits of 142 500 TNOs were integrated and their minimum semimajor axes monitored throughout the simulation. TNOs were removed from the simulation in the case of collision or ejection.
@@
Figure 4: Average fraction of cold TNOs (in blue) and hot TNOs (in red) that survived 4.54.5\,Gyr as a function of $T$ (and $I_{\infty}$).
@@
The end states of the TNO reservoirs in the $Ia$- and $ea$-planes are shown in Fig. 2 for the Tisserand parameters, T=2,1.5,1, 0, 0.5, 2, 2.5T=-2,\ -1.5,\ -1,\ 0,\ 0.5,\ 2,\ 2.5 and $2.8$ (as in Fig. 1). Hot TNOs are shown in red and cold TNOs in blue. Unstable orbits are shown in black at the last sampled epoch before instability. Also shown are the TNOs’ distributions of minimum semimajor axes achieved over 4.5 Gyr and the corresponding eccentricities and inclinations for all orbital types stable and unstable.
@@
The reservoirs’ semimajor axis extent is similar to the initial state’s. TNOs scattered by the combined effects of the giant planets, Galactic tide and passing stars are also present far outside the reservoirs’ domain. Secular perturbations cause a larger dispersion in eccentricity and inclination for hot TNOs. Cold TNOs remain mainly with low-eccentricity orbits as evidenced by their clustering along the inclination curve of circular orbits (5). As the Tisserand parameter increases toward the polar corridor boundary’s $T=2$, instability is greater and the reservoirs are depleted. Examination of the minimum semimajor axis distributions shows that starting from T=2T=-2 minimum perihelia are distributed between Saturn and Neptune but the perihelion range narrows around Neptune’s orbit outside the polar corridor as $T$ increases and initial inclinations get smaller.
@@
The Tisserand inclination pathways with respect to Neptune for the bulk of the reservoirs’ material are generally followed for all values of the Tisserand parameter. The pathways are shown in orange in the inclination panels of Fig. 2. A novel process related to Saturn adds secondary inclination pathways that are present for $T\leq 2$ and prominent inside the polar corridor. Saturn appears to retain a small fraction of reservoir material with perihelia at Saturn’s orbit whereas their orbits are pushed back outside Neptune’s. This material follows two specific Tisserand inclination pathways with respect to Saturn regardless of the initial Tisserand parameter, $T$. These are $I_{\rm Tiss}(a,-2,a_{\rm S})$ and $I_{\rm Tiss}(a,2,a_{\rm S})$ (2) where $a_{\rm S}$ is Saturn’s semimajor axis. The pathways correspond to inclinations far from Neptune $I_{\infty}=135^{\circ}$ and $I_{\infty}=45^{\circ}$ (3) and appear as the green curves in the inclination panels of Fig. 2. These secondary populations are prominent for $T=0$ and $T=0.5$ in Fig. 2 and their density decreases outside the polar corridor as $T$ tends to 2-2. Investigating Saturn’s secondary inclination pathways is beyond the scope of this paper.
@@
Figure 3 shows the ratio of the number of surviving hot TNOs to their initial number, denoted $f_{\rm TNO}$, for different initial semimajor axes as a function of the Tisserand parameter. For each semimajor axis, all the different perihelia are included in the statistic – the perihelion dependence is discussed in the next section. Four semimajor axis groups are visible. First, TNOs with initial $a$ in the range [70:100] au follow a similar distribution to that of the dynamical time in the three-body problem found in Paper IV (Fig. 7 therein). The dynamical time can be considered a proxy for $f_{\rm TNO}$. The three-body dynamical time versus Tisserand parameter, has three peaks: a prominent one near $T\sim-0.7$, a moderate one near $T\sim 0.8$ and a smaller one near $T\sim 2.2$. The $6$-body plus passing stars configuration of the present framework displaces the three peaks to $T\sim-1.5$ with $f_{\rm TNO}\sim 60\%$ for $a=80$ au, to $T\sim 0.5$ with $f_{\rm TNO}\sim 50\%$ for $a=70$ au and to $T\sim 2$ with $f_{\rm TNO}\sim 30\%$ for $a=100$ au. This group’s range of [70:100] au is reminiscent of the semimajor axis probability distributions peaks discussed in Sects. 1 and 2.
@@
The ensembles with $a=50$ au and $a=60$ au form the second group whose retrograde peak has been all but erased in favor a prominent peak at $T=0.5$ and $f_{\rm TNO}\sim 60\%$ for $a=60$ au. The third group has a semimajor axis range [110:140] au whose $f_{\rm TNO}$ decrease steadily with increasing $T$. The fourth group is made up of the ensembles of $a=40$ au, $a=200$ au and 500 au that are mainly unstable confirming the existence of two stability boundaries, an inner one between $40$ au and $50$ au and an outer one between $140$ au and $200$ au.
@@
The values of $f_{\rm TNO}$ in the range [50:140] au confirm that TNO reservoirs are long-lived in the presence of all giant planets, passing stars and the Galactic tide. In particular, the peaks at $I_{\infty}=122^{\circ}$ (T=1.5T=-1.5) and $I_{\infty}=80^{\circ}$ ($T=0.5$) single out inclinations that have the largest current populations. The average of $f_{\rm TNO}$ over all semimajor axes for hot and cold TNOs, denoted $\langle f_{\rm TNO}\rangle$, is shown a function of the Tisserand parameter in Fig. 4. It indicates that cold TNOs on average have a 4.5 Gyr dynamical time as $\langle f_{\rm TNO}\rangle>50\%$ unsurprisingly since they mainly retain low-eccentricity orbits. Hot TNOs fractions decrease with $T$ peaking at $T\sim-1.5$ with $\langle f_{\rm TNO}\rangle\sim 40\%$, $T\sim 0.5$ with $\langle f_{\rm TNO}\rangle\sim 35\%$ and $T\sim 2$ with $\langle f_{\rm TNO}\rangle\sim 22\%$.
@@ 5 Centaur injection
Figure 5: Fraction of hot TNOs that became Centaurs as a function of $T$ (and $I_{\infty}$). In the top panel, $f_{\rm cent}$, corresponds to the initial semimajor axis given next to each curve. In the bottom panel, the cold and hot TNO fractions averaged over initial semi-major axes, $\langle f_{\rm cent}\rangle$, are shown in blue and red, respectively. The black line is $\langle f_{\rm cent}\rangle(\%)=17+4.3T$.
The ratio of the number of TNOs whose semimajor axes became smaller than Neptune’s over 4.5 Gyr, to the initial TNO number is denoted $f_{\rm cent}(a,T)$. The initial TNO number of each semimajor axis includes the full perihelion range and the Centaur number includes stable and unstable orbits. The top panel of Fig. 5 shows $f_{\rm cent}(a,T)$ for hot TNOs and different semimajor axes as a function of the Tisserand parameter. The Centaur fraction increases with the Tisserand parameter at a larger rate as the semimajor axis is closer to Neptune. TNOs with low to moderate inclination are injected in greater numbers than high-inclination TNOs as their corresponding reservoirs are depleted because of instability. The hot TNO ensembles with $a=200$ au and 500 au contribute smaller Centaur numbers also because of instability. On the other hand, instability for TNOs with $a=40$ au translates into the largest injection rate because of its proximity to Neptune. For Tisserand parameters T<1T<-1, Centaur fraction is finite indicating that the presence of the giant planets and passing stars breaks Neptune’s reflecting ability that was evidenced in the three-body problem with the existence of $a_{\rm refl}$ (4) (Papers III and IV) – further evidence is given in the next section.
We defined $\langle f_{\rm cent}\rangle$ as the semimajor axis averaged Centaur fraction in the stable region. The TNO ensembles of $a=40$ au, 200 au and 500 au are excluded from the hot TNO statistics because they lie outside the stability region. This will be applied in the remainder of the paper when averaging over semimajor axes. The bottom panel of Fig. 5 shows that $\langle f_{\rm cent}\rangle$ of hot and cold TNOs are remarkably similar. The hot TNO $\langle f_{\rm cent}\rangle$ follows a simple linear expression that depends on the Tisserand parameter as follows $\langle f_{\rm cent}\rangle(\%)=17+4.3T$.
@@
More information on the injection process is found by examining the dependence of the Centaur number on perihelion for hot TNOs and on semimajor axis for cold TNOs. Figure 6 shows non-averaged $f_{\rm cent}(a,T,q)$ along with the corresponding dynamical time, $\tau_{d}(a,T,q)$, for the ensembles with an initial $a=80$ au. As stated in Sect. 2, each $(a,T,q)$ ensemble has 100 TNOs and $f_{\rm cent}(a,T,q)$ is the fraction of those among them that became Centaurs. An ensemble’s dynamical time is its median lifetime. It is meant as a stability indicator and not a dynamical time in the classical sense that is associated with a single orbit and derived from its clones’ evolution. As such, the TNOs’ dynamical time, $\tau_{d}$, shown here is an average of the classical dynamical times of the orbits that make up the $(a,T,q)$ ensemble. We find, unsurprisingly, that Centaur injection tends to increase with smaller perihelia for low to moderate inclinations ($T=2.8$). However, the perihelion dependence become shallower for reservoirs inside the polar corridor ($1\leq T\leq 2$) suggesting that it is secular perturbations and not the initial perihelion that control Centaur injection – see the panel of $T=1$. The dynamical time is an increasing function of perihelion and reaches 4.5 Gyr in the polar corridor and for T>1T>-1. Its dependence on the Tisserand parameter inside the polar corridor is unclear but could be depend on the vicinity of $T$ to the boundaries T=1T=-1 and $T=2$. Different initial semimajor axes of hot TNOs in the region [50:140] au have similar $f_{\rm cent}(a,T,q)$ and dynamical times as a function of perihelion. The median and maximum dynamical times for the different perihelia respectively are 3 Gyr and 4.5 Gyr, with the latter achieved in the polar corridor. The median and maximum dynamical times are 2.4 Gyr and 3.5 Gyr for $a=200$ au and 500 au, whereas for $a=40$ au, they are 0.48 Gyr and 1.5 Gyr.
Figure 7: Dynamical time (green) and Centaur fraction (purple) for initially cold TNOs as a function of the semimajor axis for different Tisserand parameters. The dashed blue line is $\tau_{d}=4.5$ Gyr. The empty semimajor axis ranges indicate the absence of Centaur injection.
The Centaur fraction, $f_{\rm cent}(a,T,q)$, and dynamical time of cold TNOs, $\tau_{d}(a,T,q)$, are shown in Fig. 7. The number of injected Centaurs decreases exponentially with semimajor axis so that most injected Centaurs come for the smallest semimajor axes. This is visible on the dynamical times as they reach 4.5 Gyr as soon as $a\geq 60$ au regardless of the Tisserand parameter. It is interesting that regardless of which semimajor axis and perihelion contributes most to Centaur injection, the average Centaur fractions, $\langle f_{\rm cent}\rangle$, of the hot and cold reservoirs are similar (Fig. 5 bottom panel). Understanding this similarity’s origin may reveal the common dynamical process underlying Centaur injection.